OpenStudy (anonymous):

The grades on the last math exam had a mean of 88%. Assume the population of grades on math exams is known to be normally distributed with a standard deviation of 6. What percent of students earn a score between 72% and 90%? 0.6255 0.5843 0.6754 0.5346

OpenStudy (.whitedragon.):

@ganeshie8

OpenStudy (anonymous):

@mane1

OpenStudy (anonymous):

yes

OpenStudy (.whitedragon.):

type something brooklyn ill give medal 2 u

OpenStudy (kirbykirby):

You need to find \(P(72<X<90)\). You must standardize your random variable. \[P(72<X<90)=P\left( \frac{72-88}{6}<\frac{X-88}{6}<\frac{90-88}{6}\right)=P(-2<Z<\frac{1}{3})\].

OpenStudy (anonymous):

Okay thank you! I just didn't understand how to standardize it with two numbers. So it's the same as with one?

OpenStudy (anonymous):

now I just find the z score right?

OpenStudy (kirbykirby):

oops \(P\left( \frac{-8}{3}<Z<\frac{1}{3}\right)\)

OpenStudy (kirbykirby):

i divided by 8 instead f 6 by accident

OpenStudy (anonymous):

It's fine!

OpenStudy (kirbykirby):

So yeah you look at the z-scores of -8/3 = -2.66, and z=1/3=0.33 on a normal table. And determine what the probability is between those 2 values.

OpenStudy (anonymous):

help mane1

OpenStudy (anonymous):

urgent

OpenStudy (anonymous):

Wait so how do I do that? I am actually an intelligent person I just missed a day of class and my professor didn't explain it too well to me.

OpenStudy (kirbykirby):